# Continuous Language Model Interpolation for Dynamic and Controllable Text Generation

### v1

@@ 3.1 Linear interpolation for a single attribute dimension
We first explore the effect of moving along the vector between a single pair of fine-tuned anchor models. We note that $\alpha=0$ and $\alpha=1$ correspond to the two fine-tuned anchor models, while $\alpha\in(0.0,1.0)$ is an interpolation along the vector between the two models and $\alpha\in(-\infty,0.0)\cup(1.0,\infty)$ is a linear extrapolation along the vector between the models.
@@
Linear interpolation: Figure 2 shows the effect of $\alpha$ on attribute score. For all of the attributes, when interpolating between the two fine-tuned models (Figure 1(a)), as $\alpha$ increases, there is a smooth increase in the attribute score for all of the control dimensions. The model output quality also remains high, as for every attribute the perplexity in the interpolation region is either less than or between the perplexities of the two fine-tuned models (Figure 3). These results indicate that for one control attribute, interpolating between two endpoint models yields fine-grained control over the model outputs. Furthermore, similarly to Dekoninck et al. (2024), the trend of increase with $\alpha$ appears linear in some cases (and nonlinear in others). For the majority of the attribute dimensions (politeness, formality, and simplicity) we observe a linear increase in the score as $\alpha$ increases in the interpolation region. On the other hand, the other control dimensions (sentiment and humor) have a nonlinear increase in attribute score with $\alpha$ due to plateaus at one or more of the extremes.
@@
Linear extrapolation: Figure 1(b) shows the attribute scores when extrapolating linearly beyond the two fine-tuned models along the vector between them. We find that even beyond the region of interpolation between the two fine-tuned models, there is a small stable extrapolation regime up to $\alpha$ values of around $-1$ and $2$ (Figure 1(b)). In this region, for many of the attributes, the attribute score continues to behave predictably as $\alpha$ is increased. However, beyond the stable extrapolation values, there is an unstable extrapolation regime where the attribute score changes unpredictably as $\alpha$ is varied. This is likely due to the model output quality degrading, since as shown in Figure 3, the model perplexity increases sharply starting near the edges of the stable extrapolation regime. While prior work has shown that linear weight extrapolation can be used for tasks such as model unlearning (Ilharco et al., 2023; Zhang et al., 2023), these results provide a cautionary tale against extrapolating too far, as they suggest that this ability only extends to a certain threshold before the attribute score and model outputs become unpredictable due to poor quality outputs. For the remainder of our experiments, we thus focus on the interpolation regime.
@@ 3.2.1 Parametrization of the convex hull
Fine-grained analysis of the interpolation parameter $\alpha$: We find that when interpolating across up to five attribute dimensions, modifying the weight parameters $\lambda_{i}$ and $\alpha_{i}$ results in predictable, fine-grained control over the attribute scores for the desired attributes while having a comparatively small effect on the remaining attributes. Each spider plot in Figure 4 shows that increasing the $\alpha_{i}$ parameter for interpolating between the fine-tuned models increases the attribute score for the $i$th attribute while the other scores remain fairly constant. Similarly, as the model mixture parameter $\lambda_{i}$ increases, the effect on the attribute score of changing $\alpha_{i}$ increases. While there is also more effect on the other attributes as $\lambda_{i}$ increases, this effect is still comparatively small in relation to the effect on the desired attribute.
(a) Effect of $\lambda$ on score when $\alpha_{i}=0$
(b) Effect of $\lambda$ on score when $\alpha_{i}=1$
Figure 5: Effect of $\lambda_{i}$ on the attribute scores for multiple attribute dimensions. For each control dimension $i$ and combination of parameters from Figure 4 where $\alpha_{i}=0$ (4(a)) and Figure 11 where $\alpha_{i}=1$ (4(b)), we report the attribute score versus $\lambda_{i}$. In 4(a), as $\lambda_{i}$ increases, the attribute score decreases smoothly toward the score for the $\alpha_{i}=0$ model, and in 4(b), as $\lambda_{i}$ increases, the attribute score increases smoothly toward the $\alpha_{i}=1$ score.
Fine-grained analysis of the mixing parameter $\lambda$: The effect of $\lambda$ is further demonstrated in Figure 4(a), which plots the scores from Figure 4 for each attribute dimension and $\lambda_{i}$ value when $\alpha_{i}=0$ across all values of $\lambda_{j}$ for the other control dimensions. We note that in this case, increasing $\lambda_{i}$ should upweight the $\alpha_{i}=0$ model and thus decrease the attribute score. Figure 4(b) is the analogous plot for $\alpha_{i}=1$, showing the scores averaged across all combinations of weights from Figure 11 when $\alpha_{i}=1$ for the attribute dimension being plotted. In this case, increasing $\lambda_{i}$ should increase the score, since the weight of the $\alpha_{i}=1$ model is increasing. Combined, these plots shows that as $\lambda_{i}$ increases, the output scores move smoothly toward the desired extreme model for both the $\alpha_{i}=0$ and $\alpha_{i}=1$ case, showing that the $\lambda$ parameter also provides fine-grained control over the model outputs.
Figure 6: Effect of $\lambda_{i}$ on interpolation between the sentiment, politeness, and simplicity dimensions for $\alpha_{i}=1$. The vertices of the triangle represent the models with $\alpha_{i}=1$ for each of the three attribute dimensions. The scores in the simplex of $\lambda$ weights between the three control dimensions smoothly interpolate between the three extreme models.
@@
Changing mixing parameters $\lambda$ for multiple attributes at once: We also analyze the relationship throughout the whole simplex of $\lambda$ weights for sets of three control dimensions in Figures 6 and 7 (as well as Figures 12-21 in the Appendix). For each set of three attributes listed, these plots show the scores in the three dimensional simplex of mixing weights $\lambda$ for which $\sum_{i}\lambda_{i}=1$. The value of the interpolation weight $\alpha_{i}$ for each of the attributes is equal to 1 in Figures 6 and 7, so increasing the $\lambda$ weight of each attribute should increase the attribute score. We find that surprisingly, there is very limited entanglement between the majority of the combinations of attributes (such as in Figure 6). In these cases, we observe an approximately even increase in score as $\lambda_{i}$ for a given attribute dimension increases, regardless of the other $\lambda_{j}$ parameters.
However, in the case of humor in the humor-formality-simplicity simplex and formality in the humor-formality-simplicity simplex with $\alpha_{i}=1$ (Figure 7) and the sentiment-politeness-formality simplex with $\alpha_{i}=1$ (Figure 17), we observe regions at the corners of the simplex that are close to the other fine-tuned models and have a high attribute score. This is because these other models are correlated with a positive attribute score, so the mixture of models is the most neutral model. Nevertheless, this still has a limited effect on the attribute score, since even in these cases with correlations, the score still has the expected behavior unless the mixing weight $\lambda_{j}$ is greater than around $0.4$ to $0.6$ for the correlated control dimensions. This indicates that in practice, the model has smoothly increasing attribute scores with $\lambda_{i}$ for all pairs of attributes when $\lambda_{j}$ for the other attribute dimensions remains sufficiently low.
@@
3.2.2 Fine-tuned models and correlations
Given the results from the simplex plots, we analyze the relationships between the fine-tuned endpoint models to better understand the attribute score correlations. Figure 8, which plots the average cosine similarity between the LoRA layers of each pair of models, shows that the LoRA weights are relatively orthogonal to each other in most cases. We hypothesize that the lower orthogonality between each pair of endpoint models for the same attribute is because the models are trained on similar datasets. This is supported by the fact that the simple and complex models are the most orthogonal of the pairs of endpoint models and they are the only two models trained on different datasets rather than different classes from the same dataset. In addition, the humor models tend to deviate the most from orthogonality with the other models (such as politeness), so this may provide a partial explanation for why some of the other models were correlated with a higher humor score.
Figure 8: Cosine similarity of LoRA weights averaged across layers between each pair of fine-tuned models. The LoRA weights are all relatively orthogonal to each other, except some of the two endpoint models for the same attribute are less orthogonal to each other, as well as the politeness and humorous models.
@@ 4.1 Controllable text generation (CTG)
As it is crucial to constrain generated text in many downstream applications, CTG has been a recent focus of NLP research. Methods such as CTRL (Keskar et al., 2019) and GeDI (Krause et al., 2021) pretrain language models on text prepended with control codes and generate text conditioned on the desired control. However, these methods require pretraining a new model if new controls are added, which is computationally expensive. To mitigate these issues, a variety of methods have been proposed to perform CTG without additional language model training. For example,  Khalifa et al. (2021); Pascual et al. (2021); Yang & Klein (2021); Dekoninck et al. (2024) constrain language model outputs by modifying their output probability distributions.  Li & Liang (2021); Qian et al. (2022) learn prefixes and Han et al. (2023) train a linear factor in the word embedding space.  Subramani et al. (2022); Hernandez et al. (2023); Li et al. (2023); Turner et al. (2023) control model outputs by changing activations at inference time.  Zhou et al. (2023) use instruction tuning for CTG.
In this prior CTG research, only  Dekoninck et al. (2024) show that their method is composable and achieves fine-grained control over multiple attributes at once. However, as this method requires composing multiple models at inference time, the inference cost is significantly higher than inference from a single weight-interpolated model, especially as the model size and number of controlled attributes increases. In addition, combining low-rank fine-tuning weights instead of probability distributions allows for more complex relationships between the models to be taken into account when composing them, which will likely allow for greater flexibility as the number of controlled attributes increases.
@@ 4.2 Weight interpolation
Our work builds on prior work on linear weight interpolation, such as task vectors (Ilharco et al., 2023), parameter-efficient task vectors (Zhang et al., 2023), and model souping (Wortsman et al., 2022), as we use linear interpolation and weighted model averaging as the basis for our analysis. Prior work in this domain has focused mainly on improving multitask performance when composing fully fine-tuned models (Matena & Raffel, 2021; Yadav et al., 2023; Ortiz-Jimenez et al., 2023; Ramé et al., 2023) or parameter-efficient fine-tuned models (Huang et al., 2024; Jiang et al., 2024). However, these methods all differ from our work, since they focus on combining model weights to improve a single multitask objective rather than analyzing performance across a wide range of flexible, diverse objectives. These approaches are orthogonal to our work and could be used in conjunction with it to better combine the $\alpha$-interpolated models. Perhaps most similar to our work are methods that interpolate between the weights of fine-tuned models to control over a range of outputs (Gandikota et al., 2023; Nylund et al., 2023). However, Gandikota et al. (2023) focus on the vision domain and use a fine-tuning objective specific to diffusion models, and Nylund et al. (2023) only analyze control over the time dimension.
5 Conclusion and future work
@@
In this work, we show that continuous linear interpolation between low-rank fine-tuned models can be used to parametrize the models in their convex hull. We achieve fine-grained, predictable control over multiple attributes of style at once by changing the interpolation weights between two anchor fine-tuned models and the mixing weights between different interpolated attribute models. We find that the interpolation profiles between models are smooth and there is surprisingly little entanglement between the models for different control dimensions. In other words, changing the weight for one attribute has a very small effect on the scores for other attributes, especially for sufficiently small mixing weights. As a result, we show that linear weight interpolation can be used to dynamically adjust to diverse sets of changing preferences and generate text that adheres to multiple controls simultaneously.
@@
Limitations and future work: The main limitation of our work is that some pairs of attributes are correlated, so when a correlated model has a large mixing weight, it can unpredictably affect other control attributes. Thus, a natural extension of this work would be to investigate whether this correlation is inherent to the pair of tasks or if it can be eliminated. For example, text that is more polite might always be more formal. However, it may be the case that some correlations can be reduced by regularizing the LoRA updates to be more orthogonal to each other or by merging the $\alpha$-interpolated using more sophisticated methods that have recently shown improvement over naive weight averaging in the multitask setting (Matena & Raffel, 2021; Yadav et al., 2023; Ortiz-Jimenez et al., 2023; Ramé et al., 2023).
Another potential focus of future work could be to extend the extrapolation results to multiple control dimensions to analyze whether it is possible to reliably generate text beyond the fine-tuned models when controlling multiple attributes at once. This could be useful to further extend the range of control over the model outputs.
6 Ethics Statement

### v2

@@ 3.1 Linear interpolation for a single attribute dimension
We first explore the effect of moving along the vector between a single pair of fine-tuned anchor models. We note that $\alpha=0$ and $\alpha=1$ correspond to the two fine-tuned anchor models, while $\alpha\in(0.0,1.0)$ is an interpolation and $\alpha\in(-\infty,0.0)\cup(1.0,\infty)$ is an extrapolation along the vector between the models.
Linear interpolation: Figure 2 shows the effect of $\alpha$ on attribute score when interpolating between each pair of fine-tuned anchor models. As $\alpha$ increases, there is a smooth and predictable increase in the attribute score for all of the control dimensions. Furthermore, linear interpolation follows the trend of the ground truth fine-tuned models more closely than either of the baselines while achieving similar perplexity and better diversity (Table 1). This demonstrates that interpolation can be used to approximate this continuous class of intermediate fine-tuned models by fine-tuning only $2$ endpoint models. In contrast, for the model arithmetic baseline, the unpredictability of the attribute score in the politeness and humor dimensions suggests that prompt-conditioned models provide somewhat inconsistent control in the continuous interpolation setting in comparison to fine-tuned models. We find similar results for other language models (Appendix A.7) and diversity metrics (Table 6).
@@
These results indicate that for one control attribute, weight interpolation between two endpoint models yields fine-grained control over the model outputs. Furthermore, the trend of increase with $\alpha$ appears linear in some cases. For the majority of the attribute dimensions (politeness, formality, and simplicity) we observe a linear increase in the score as $\alpha$ increases in the interpolation region. On the other hand, the other control dimensions (sentiment and humor) have a nonlinear increase in attribute score with $\alpha$ due to plateaus at the extremes.
Figure 3: Effect of linear weight extrapolation for a single attribute dimension. For each style attribute, we report the attribute score when linearly extrapolating beyond the fine-tuned models ($\alpha<0$ and $\alpha>1$). There is a stable region where the score changes smoothly until a certain point (around $\alpha$ equal to 1-1 and 22), where performance degrades and the extrapolation is unstable.
Figure 4: Wikitext perplexity of linearly interpolated and extrapolated models. We report the average perplexity (lower is better) of each model from Figure 4 on the Wikitext test set. For the extrapolated models ($\alpha<0.0$ and $\alpha>1.0$), the perplexity increases rapidly beyond $\alpha$ values of around 1-1 and $2$. We clip the y-axis at 7.0 for readability (full plot in Figure 10).
Figure 5: Effect of $\alpha_{i}$ and $\lambda_{i}$ on 5-dimensional interpolation. For each attribute, we show the attribute scores for models with the given $\alpha_{i}$ and $\lambda_{i}$ parameters, with all four other $\alpha_{j}=0$ and $\lambda_{j}=(1-\lambda_{i})/4$. We find that increasing $\alpha_{i}$ consistently increases the attribute score and increasing $\lambda_{i}$ consistently increases the effect of $\alpha_{i}$.
@@
Linear extrapolation: Figure 4 shows the attribute scores when extrapolating linearly beyond the two fine-tuned models along the vector between them. We find that even beyond the region of interpolation between the two fine-tuned models, there is a small stable extrapolation regime up to $\alpha$ values of around 1-1 and $2$ (Figure 4). In this region, for many of the attributes, the attribute score continues to behave predictably as $\alpha$ is increased. However, beyond the stable extrapolation values, there is an unstable extrapolation regime where the attribute score changes unpredictably as $\alpha$ is varied. This is likely due to the model output quality degrading, since as shown in Figure 4 and Figure 11, the model perplexity increases sharply and the diversity decreases starting near the edges of the stable extrapolation regime. While prior work has shown that linear weight extrapolation can be used for tasks such as model unlearning (Ilharco et al., 2023; Zhang et al., 2023), these results provide a cautionary tale against extrapolating too far, as they suggest that this ability only extends to a certain threshold before the attribute score and model outputs become unpredictable due to poor quality outputs. For the remainder of our experiments, we thus focus on the interpolation regime.
@@ 3.2.1 Parametrization of the convex hull
Analysis of interpolation parameter $\alpha$ and $\lambda$: We find that when interpolating across up to five attribute dimensions, modifying the weight parameters $\lambda_{i}$ and $\alpha_{i}$ results in predictable, fine-grained control over the attribute scores for the desired attributes while having a comparatively small effect on the remaining attributes. Each attribute plot in Figure 5 shows that increasing the $\alpha_{i}$ parameter for interpolating between the fine-tuned models increases the attribute score for the $i$th attribute in a predictable manner. Similarly, as the model mixture parameter $\lambda_{i}$ increases, the effect on the attribute score of changing $\alpha_{i}$ increases.
Figure 6: Effect of $\lambda_{i}$ on interpolation between the sentiment, politeness, and simplicity dimensions for $\alpha_{i}=1$. The vertices of the triangle represent the models with $\alpha_{i}=1$ for each of the three attribute dimensions. The scores in the simplex of $\lambda$ weights between the three control dimensions smoothly interpolate between the extreme models.
@@
Changing mixing parameters $\lambda$ for multiple attributes at once: We also analyze the relationship throughout the whole simplex of $\lambda$ weights for sets of three control dimensions in Figures 6 and 7 (as well as Appendix Figures 29-42). For each set of three attributes listed, these plots show the scores in the three dimensional simplex of mixing weights $\lambda$ for which $\sum_{i}\lambda_{i}=1$. The value of the interpolation weight $\alpha_{i}$ for each of the attributes is equal to 1 in Figures 6 and 7, so increasing the $\lambda$ weight of each attribute should increase the attribute score. We find that surprisingly, there is very limited entanglement between the majority of the combinations of attributes (such as in Figure 6), likely because their LoRA weights are relatively orthogonal to each other (Appendix Figure 44). In these cases, we observe an approximately even increase in score as $\lambda_{i}$ for a given attribute dimension increases, regardless of the other $\lambda_{j}$ parameters.
However, in some cases, such as humor in the humor-formality-simplicity simplex and formality in the humor-formality-simplicity simplex with $\alpha_{i}=1$ (Figure 7), we observe regions at the corners of the simplex that are close to the other fine-tuned models and have a high attribute score. This is because these other models are correlated with a positive attribute score, so the mixture of models is the most neutral model. Nevertheless, this still has a limited effect on the attribute score, since even in these cases with correlations, the score still has the expected behavior unless the mixing weight $\lambda_{j}$ is greater than around $0.4$ to $0.6$ for the correlated control dimensions. This indicates that in practice, the model has smoothly increasing attribute scores with $\lambda_{i}$ for all pairs of attributes when $\lambda_{j}$ for the other attribute dimensions remains sufficiently low.
@@
Figure 8: Weight interpolation has entanglement lower than or comparable to prior approaches. For each pair of dimensions, we fix (scaled) $\alpha_{i}=1$ for the dimension in each row and vary (scaled) $\alpha_{j}$ between 0 and $1$ for the dimension in each column. We set $\lambda_{i}=\lambda_{j}=0.5$. Then, we report entanglement as the area under the curve of absolute value of change in attribute score as $\alpha_{j}$ increases. Weight interpolation is less entangled than DExperts (Liu et al., 2021) and has similar entanglement to model arithmetic (Dekoninck et al., 2024). These results hold across different language model sizes and families (Appendix A.9).
3.2.2 Entanglement analysis
Given the results from the simplex plots, we analyze the entanglement between each pair of dimensions to better understand the attribute score correlations. For every pair of dimensions $(i,j)$, if the two attributes are not entangled, it should be the case that if $\alpha_{i}$ is held constant, then changing $\alpha_{j}$ does not affect the attribute score. We can thus evaluate the amount of entanglement for a given $\alpha_{j}$ value by computing the absolute value of the difference between the attribute score for that $\alpha_{j}$ value and the attribute score when $\alpha_{j}=0$. We do so for scaled $\alpha_{i}=1$ (for $\alpha_{i}=0$ see Appendix Figure 20 and for additional language models see Appendix A.9) and $\alpha_{j}\in\{0,0.25,0.5,0.75,1.0\}$ with $\lambda_{i}=\lambda_{j}=0.5$. We then compute the area under the curve (AUC) of the entanglements for these $\alpha_{j}$ values in Figure 8. We find that the entanglement for weight interpolation is comparable to that of model arithmetic (Dekoninck et al., 2024), but less than that of DExperts (Liu et al., 2021).
@@ 4.1 Controllable text generation (CTG)
As it is crucial to constrain generated text in many downstream applications, CTG has been a recent focus of NLP research. Methods such as CTRL (Keskar et al., 2019) and GeDI (Krause et al., 2021) pretrain language models on text prepended with control codes and generate text conditioned on the desired control. However, these methods require pretraining a new model if new controls are added, which is computationally expensive. To mitigate these issues, a variety of methods have been proposed to perform CTG without additional language model training. For example,  Liu et al. (2021); Khalifa et al. (2021); Pascual et al. (2021); Yang & Klein (2021); Dekoninck et al. (2024) constrain language model outputs by modifying their output probability distributions.  Li & Liang (2021); Qian et al. (2022) learn prefixes and Dathathri et al. (2019); Han et al. (2023) train additional classifiers to guide generation.  Subramani et al. (2022); Hernandez et al. (2023); Li et al. (2023); Turner et al. (2023) control model outputs by changing activations at inference time.  Kumar et al. (2021) optimize the inference decoding.  Mireshghallah et al. (2022); Qin et al. (2022) use energy-based constrained generation and  Zhou et al. (2023) use instruction tuning for CTG.
However, we emphasize that the goal of our work is not to output text with the greatest or least possible amount of a given style attribute, but instead to show that interpolation can be used to dynamically and controllably increase or decrease the amount of a given style attribute. While most of these existing methods can be used to control text for multiple style attributes, the majority of them focus on binary or multi-class control (for example controlling text so that it has positive or negative sentiment) and would need to be re-run for each combination of attribute levels, which is not feasible for "continuous" adaptive control.
As a result, among existing methods only DExperts (Liu et al., 2021) and model arithmetic (Dekoninck et al., 2024) are composable and achieve fine-grained control over multiple attributes at once. Both weight interpolation and DExperts require fine-tuning two endpoint models for each attribute, while model arithmetic uses prompt-conditioned models. However, both DExperts and model arithmetic compose multiple models at inference time, so the inference cost is significantly higher than weight interpolation, especially as the model size and number of controlled attributes increases. In addition, our experiments show that weight interpolation provides more precise fine-grained control and less entanglement than these approaches over the range of intermediate models.
@@ 4.2 Weight interpolation
Our work builds on prior work on linear weight interpolation, such as task vectors (Ilharco et al., 2023), parameter-efficient task vectors (Zhang et al., 2023), and model souping (Wortsman et al., 2022), as we use linear interpolation and weighted model averaging as the basis for our analysis. Prior work in this domain has focused mainly on improving multitask performance when composing fully fine-tuned models (Matena & Raffel, 2021; Yadav et al., 2023; Ortiz-Jimenez et al., 2023) or parameter-efficient fine-tuned models (Huang et al., 2024; Jiang et al., 2024). However, these methods all differ from our work, since they focus on combining model weights to improve a single multitask objective rather than analyzing performance across a wide range of flexible, diverse objectives. These approaches are orthogonal to our work and could be used in conjunction with it to better combine the $\alpha$-interpolated models.
Beyond multitask performance, a variety of reinforcement learning approaches have been proposed for language model alignment to user preferences. Specifically, rewarded soups first fine-tunes a model for each preference and computes the combination of interpolation parameters that maximizes the user’s reward on a validation set (Ramé et al., 2023), personalized soups uses a multi-objective reinforcement learning objective combined with parameter merging (Jang et al., 2023), and conditional language policy performs multi-objective fine-tuning to learn a set of parameters that is steerable by conditioning on the user reward at inference time (Wang et al., 2024). These methods show that reinforcement learning objectives can be combined with weight interpolation in order to optimize over a wide range of user preferences. However, they do not analyze the behavior across the full interpolation region and instead center around developing methods that obtain close to pareto-optimal reward or are more preferential to users with specified reward functions.
Perhaps most similar to our work are methods that analyze the interpolation regime between the weights of fine-tuned models over a range of outputs (Gandikota et al., 2023; Nylund et al., 2023). However, Gandikota et al. (2023) focus on the vision domain and use a fine-tuning objective specific to diffusion models, and Nylund et al. (2023) only analyze control over the time dimension.
5 Conclusion
@@
In this work, we show that continuous linear interpolation between low-rank fine-tuned models can be used to parametrize the models in their convex hull. We achieve fine-grained, predictable control over multiple attributes of style at once by changing the interpolation weights between two anchor fine-tuned models and the mixing weights between different interpolated attribute models. We find that the interpolation profiles between models are smooth and there is surprisingly little entanglement between the models for different control dimensions. In other words, changing the weight for one attribute has a very small effect on the scores for other attributes, especially for sufficiently small mixing weights. As a result, linear weight interpolation produces predictable behavior in the entire continuous space of models between the fine-tuned anchors.
Future work
@@
The main limitation of our work is that some pairs of attributes are correlated, so when a correlated model has a large mixing weight, it can unpredictably affect other control attributes. While there is lower or comparable correlation in weight interpolation as compared to other approaches, in the future it would be valuable to investigate whether this correlation is inherent to the pair of tasks or if it can be eliminated. For example, text that is more polite might always be more formal. However, it may be the case that some correlations can be reduced by regularizing the LoRA updates to be more orthogonal to each other or by merging the $\alpha$-interpolated using more sophisticated methods that have recently shown improvement over naive weight averaging in the multitask setting (Matena & Raffel, 2021; Yadav et al., 2023; Ortiz-Jimenez et al., 2023). Similarly to prior work on task vectors (Ilharco et al., 2023), correlations could also potentially be used to combine existing control attributes to make new ones (such as by combining irony and humor to control for sarcasm).
Another limitation is that the average generation attribute scores are limited to the range between the attribute scores of the fine-tuned anchor models. The single attribute extrapolation results could be expanded upon to better understand when extrapolation can be used to extend the range of the control attribute style. We also only consider controlling for text style in our experiments, but analyzing linear interpolation when controlling for other attributes such as topics is a possible direction for future work.
Acknowledgments
SK is supported by the National Science Foundation Graduate Research Fellowship under Grant No. DGE 2140743. DAM and SK acknowledge support from the Kempner Institute, the Aramont Fellowship Fund, and the FAS Dean’s Competitive Fund for Promising Scholarship.
Broader Impacts